The amplitude of a pendulum goes from 32 degrees from vertical to 24 degrees from vertical over the course of 20 seconds. Its gradual loss of energy may be adequately modeled as damped SHM. After another 40 seconds, what is the amplitude of the pendulum?
Answer #1

0=0 24 = 32 e bx20 now afton another 40 sec means al = 40+20= 60 sec. 0= do esta 0=32 e ble 0 = 32 (e ) 0=32 33 0 - 32 × 27 3